OpenStudy (the.lit.owl_23):

please help y = 2/5x - 3 graphing

OpenStudy (the.lit.owl_23):

(sigh)

OpenStudy (noname):

\[y=\frac{2}{5x-3}\] like that?

OpenStudy (the.lit.owl_23):

I think so thx dude

OpenStudy (noname):

It looks like a transformation of the parent function 1/x. A vertical scale change of 2 Horizontal scale change of 5 Horizontal change of 3 to the right.

OpenStudy (royalranger):

Or is it: \[\frac{ 2 }{ 5 }x -3\]

OpenStudy (the.lit.owl_23):

yeaaaaaaaaaaaaaaaa

OpenStudy (the.lit.owl_23):

y = 2/5 -3

OpenStudy (royalranger):

y = mx + b What is your slope and what is your y-intercept?

OpenStudy (noname):

oh, so my equation was wrong?

OpenStudy (the.lit.owl_23):

no

OpenStudy (royalranger):

Alright so which equation is it? you need to be clear.

OpenStudy (troy121101):

@royalranger is correct, 2/5 would be your slope, and -3 your y-intercept

OpenStudy (troy121101):

so one of the points you need to graph is the '-3' on the y-axis, because that is where the y intercepts the origin line on the graph

OpenStudy (troy121101):

and then you need to figure out the x-intercept to find the slope

OpenStudy (the.lit.owl_23):

(x,y)

OpenStudy (troy121101):

right you you should have (0,-3) graphed on your sheet

OpenStudy (the.lit.owl_23):

so far I have (0,-3)

OpenStudy (the.lit.owl_23):

yep

OpenStudy (troy121101):

good, now, a way to find this out, is since the numbers in "2/5" are positive, you would go up two, and right 5, and there is your x-intercept and slope form

OpenStudy (the.lit.owl_23):

um

OpenStudy (troy121101):

or up 2 and left 5, starting from your y-intercept graph point

OpenStudy (the.lit.owl_23):

7?

OpenStudy (troy121101):

or (0, -3), starting from that, you move up 2, and left or right 5, I am not sure which

OpenStudy (the.lit.owl_23):

the graph is not right in front of me :(

OpenStudy (troy121101):

okay, give me a second...

OpenStudy (the.lit.owl_23):

would it be (7,-3)

OpenStudy (royalranger):



OpenStudy (troy121101):

@royalranger is correct there

OpenStudy (the.lit.owl_23):

thx soo it is (5,-3)

OpenStudy (the.lit.owl_23):

??????? guys :(

OpenStudy (royalranger):

Do you know what y = mx + b means?

OpenStudy (the.lit.owl_23):

I am sorry I am not very good a graphing

OpenStudy (the.lit.owl_23):

yes

OpenStudy (the.lit.owl_23):

Slope-intercept form: y = mx + b

OpenStudy (troy121101):

Okay, so we can substitute 'm' for 2/5, as it being your slope, and 'b' for -3 as it being your y-intercept... Like this: \[y = \frac{ 2 }{ 5 }x + -3\]

OpenStudy (the.lit.owl_23):

ok

OpenStudy (troy121101):

oh boy, now I am getting confused, lol

OpenStudy (troy121101):

@royalranger can you add on to this?